OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (5, 6) and (9, 14)? y=___x+ ___

OpenStudy (susanna):

find the slope first

OpenStudy (anonymous):

(y2- y1)/(x2-x1) right?

OpenStudy (susanna):

\[\frac{ 14-6 }{ 9-5 }=\frac{ 8 }{ 4 }=2\]

OpenStudy (anonymous):

whale then.

OpenStudy (susanna):

use y-y=m(x-x) now

OpenStudy (susanna):

\[y-6=2(x-5)\]

OpenStudy (susanna):

I used (5,6)

OpenStudy (anonymous):

2x-4

OpenStudy (susanna):

yes y=2x-4